# One-to-one correspondence reconstruction at the electron-positron Higgs factory

### v1

@@ Abstract
Figure 5: Top left: Neutral hadron energy resolution using hadronic calorimeter (HCAL, blue and green lines) or TOF (red and yellow lines). Top right: Jet origin identification performance (characterized by flavor tagging efficiencies and charge flip rates) in two scenarios: Scenario 1 (S1) assumes perfect identification of leptons and charged hadrons, while Scenario 2 (S2) extends S1 by further assuming perfect identification of $K^{0}_{L}/K^{0}_{S}$, $n/\bar{n}$, and Λ/$\Lambda/\bar{\Lambda}$. Bottom: Anticipated upper limits on branching ratios of Higgs exotic and rare decays with different performances of JOI and BMR.
Since 1-1 correspondence reconstruction provides the particle mass information via PID, the energy of visible particles can be determined by the TOF. 5D calorimetry provides TOF measurements for both charged and neutral particles. For charged particles, their energies/momenta can be determined by three methods: shower energy via calorimeter, track momentum via tracker, and TOF via calorimeter, while the TOF method has a comparative advantage in the forward region. For neutral particles, their energies can be determined by both the shower energy and the TOF via calorimeter, with TOF showing a significant advantage for low-energy neutral hadrons, as demonstrated in the top left panel of Fig. 5. Therefore, by combining these measurements of particle energy, such as selecting the most suitable method according to the particle species and their kinematics, we can significantly improve the energy/momentum resolution of different particles, and consequently improve the BMR. Considering the improvement in particle energy measurement can also benefit pattern recognition and confusion control, we estimate that this approach could further enhance the BMR by 15–20%, meaning that a BMR of 2.2–2.4% could, in principle, be achieved.
1-1 correspondence reconstruction can also significantly enhance the JOI performance. JOI is the procedure to determine the type of quark/gluon (11 types are typically considered: $b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, $g$) from which a jet originates. In a recent study [36], we realize JOI with the CEPC CDR baseline detector. It can simultaneously identify $b$, $c$, $s$ quarks with efficiencies of 70–90% and $u$, $d$ quarks with efficiencies of 40%, meanwhile, the charge flip rates of quarks and anti-quarks are controlled to be 10–20%, with an ideal lepton and charged hadron identification. JOI significantly boosts the accuracy of flavor-sensitive physics measurements, such as $H\to s\bar{s}$ and $H\to sb$ decays [36], $|V_{cb}|$ [41], $\phi_{s}$ [42], and weak mixing angle [43] measurements. 1-1 correspondence provides the identification of $K^{0}_{L}$, $n$, and $\bar{n}$, while in principle the $K^{0}_{S}$ and $\Lambda$ could also be identified [44], therefore, providing the PID information of all these neutral hadrons, we observe a significant improvement on JOI performance especially the identification performance of $u/d/s$ quarks.
BMR and JOI are key performances of event reconstruction at electron-positron Higgs factories. Their improvements will benefit most physics measurements with hadronic final states. For example, the Higgs to invisible final states is a key portal to detect dark matter at Higgs factories. The original upper limit at 95% confidence level is quantified to be $\mathcal{O}(0.1\%)$ [45] with a BMR of 3.7%. Improving the BMR to 2.75%, or even 2.2%, will consequently enhance the upper limit by 10–12% [46]. Higgs exotic and rare hadronic decays are also sensitive probes to new physics. Using JOI technology, these decay modes can be typically limited to $10^{-4}$$10^{-3}$ using $\nu\bar{\nu}H$ and $\ell^{-}\ell^{+}H$ processes [36]. With the improvement in JOI alone, the constraints could be enhanced by 5–40%. Considering also the improved BMR, the sensitivities are expected to be further enhanced by another 5–10%, see the bottom panel of Fig. 5. In short, 1-1 correspondence could significantly boost the discovery power of electron-positron Higgs factories.
Based on this study, we are confident that 1-1 correspondence reconstruction will be a new paradigm for event reconstruction at the high-energy frontier. For the electron-positron Higgs factory, 1-1 correspondence reconstruction is a goal that should be pursued, as it can significantly enhance the physics reach and discovery power. It is also a goal that could be achieved—presumably very soon—through integrating innovative detector technologies, advanced reconstruction algorithms, and artificial intelligence.
@@ Acknowledgements
We would like to express our sincere gratitude to Jianchun Wang for his continuous support throughout this work. We also thank Hengyu Wang for his study on HCAL resolution, which provided an important reference for our research. This work is supported in part by the National Natural Science Foundation of China under grant No. 12042507, the National Key R&D Program of China under Contracts No. 2022YFE0116900, and the National Key Program for S&T Research and Development (2023YFA1606300). Additional support is provided by the Fundamental Research Funds for the Central Universities, Peking University.

### v4

@@ 3 Prospects
Figure 5: Top left: Neutral hadron energy resolution using calorimeter with direct energy sum of each hit [42, 50] (blue line, derived by the full detector simulation) or TOF measurement with time resolution assumed to be 10 ps at particle level (red and yellow lines). Top right: Jet origin identification performance (characterized by flavor tagging efficiencies and charge flip rates) in two scenarios: Scenario 1 (S1) assumes perfect identification of leptons and charged hadrons, while Scenario 2 (S2) extends S1 by further assuming perfect identification of $K^{0}_{L}/K^{0}_{S}$, $n/\bar{n}$, and Λ/$\Lambda/\bar{\Lambda}$. Bottom: Anticipated upper limits on branching ratios of Higgs exotic and rare decays with different performances of JOI and BMR.
4 Summary
Regarding the entire detector signal recording and event reconstruction as the process to establish a mapping between visible and reconstructed particles, 1-1 correspondence represents the ultimate goal of an isomorphic mapping. It characterizes events with observables associated with individually reconstructed particles, enabling direct and holistic signal selection algorithms that can significantly enhance the discovery power. This study demonstrates the feasibility of 1-1 correspondence reconstruction at the electron-positron Higgs factory. By employing innovative detector designs, such as high-granularity 5D calorimetry, alongside advanced algorithms like PFA and machine learning, it is possible to efficiently suppress and identify fake particles—the leading source of confusion that violates the 1-1 correspondence relationship, leading to a 25% improvement in the invariant mass resolution of hadronically decayed Higgs bosons (i.e. BMR). Using the same machine learning model, nine types of particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$, $\gamma$, $K_{L}^{0}$, $n$, $\bar{n}$) can be simultaneously distinguished with high efficiencies, greatly expanding and unifying traditional PID methods. 1-1 correspondence reconstruction also shifts the bottlenecks in experiment design, event reconstruction, and physics measurements. This study shows that the primary bottleneck in particle flow reconstruction, namely confusion, can be reduced by one order of magnitude, while the current limitations shift to detector acceptance and neutral particle identification. This study also urges the necessity of high-precision MC tools, which require profound understanding of hadronization processes, particle-matter interactions, detector responses, etc. The study also indicates that 1-1 correspondence represents an emerging paradigm for event reconstruction at the high-energy frontier. Achieving this paradigm relies on state-of-the-art detector design, innovative reconstruction algorithms, and artificial intelligence, potentially leading to a significantly deeper understanding of physics events and the underlying physics principles.
CRediT authorship contribution statement
Yuexin Wang: Investigation, Methodology, Data curation, Formal analysis, Visualization, Writing - original draft, Writing – review and editing. Hao Liang: Investigation, Methodology, Formal analysis, Visualization, Writing – review and editing. Yongfeng Zhu: Investigation, Methodology, Formal analysis, Visualization, Writing – review and editing. Yuzhi Che: Investigation, Methodology, Writing – review and editing. Xin Xia: Data curation, Validation, Writing – review and editing. Huilin Qu: Software, Writing – review and editing. Chen Zhou: Writing – review and editing. Xuai Zhuang: Writing – review and editing. Manqi Ruan: Conceptualization, Methodology, Writing - original draft, Writing – review and editing, Supervision, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgements
We would like to express our sincere gratitude to Jianchun Wang for his continuous support throughout this work. In addition, we thank Hengyu Wang for his study on HCAL resolution and thank Jianfeng Jiang for his study on event generator, which provided important references for our research. We are also grateful to Yuxuan Zhang for his review of the manuscript and valuable suggestions. This work is supported in part by the National Natural Science Foundation of China under grant No. 12042507, the National Key R&D Program of China under Contracts No. 2022YFE0116900, and the National Key Program for S&T Research and Development (2023YFA1606300). Additional support is provided by the Fundamental Research Funds for the Central Universities, Peking University.
Data availability
The code used to produce the samples and perform the analysis is available at https://code.ihep.ac.cn/wangyuexin/1-1-correspondence-reconstruction.git
Appendix A Appendix
A.1 AURORA detector
AURORA, the conceptual detector used in this study, is a PFA-oriented detector design developed from the CEPC CDR baseline [9, 51]. It has an outer radius of 5.2 m, a length of 10.5 m, and a total weight of approximately 5000 tons, with over 65% of the weight contributed by the Yoke and nearly 25% by the HCAL. Fig. 6 displays two side views of AURORA, and specific geometry parameters are summarized in Table 1. In general, AURORA features a high-granularity calorimetry with both ECAL and HCAL, a high-precision tracker system with a low material budget, a high-precision vertex detector, and a large solenoid that encloses ECAL and HCAL. Compared to the CDR baseline, the major change is the HCAL, where we replace the Glass Resistive Plate Chamber (GRPC) digital HCAL [9] with the Glass-Scintillator HCAL [39] and increase the HCAL thickness from 5 to 6 nuclear interaction lengths ($\lambda$) to reduce the longitudinal leakage. In addition, for both ECAL and HCAL, we assume each calorimeter cell can provide a time resolution of 100 ps, which is used for the PID and PFA confusion identification.
Figure 6: The $r$-$z$ (left) and $r$-$\phi$ (right) views of the AURORA detector geometry.
Table 1: AURORA detector geometry parameters.
Sub-detector Thickness (mm) Inner radius (mm) Outer radius (mm) Length (mm) Volume (m3) Transverse cell size #Layers #Channels
Vertex - - 16–60 125–250 - $25\times 25$ $\upmu$m2 6 $5.3\times 10^{8}$
Si-strip Tracker - - 155 300 1810 736 1288 4600 - 20 $\upmu$m $\times$ 2 cm 3 $3.0\times 10^{7}$
TPC - 300 1800 4700 47 $1\times 6$ mm2 220 $2.9\times 10^{6}$
ECAL 173 1845 2018 5250 15 $1\times 1$ cm2 30 $2.5\times 10^{7}$
HCAL 1145 2072 3250 7590 180 $2\times 2$ cm2 48 $1.8\times 10^{7}$
Solenoid 700 3275 3975 7750 120 - - -
Yoke 1200 4000 5200 10500 470 - - -
A.2 Monte Carlo simulation and reconstruction
A total of 1 million $ZH\to\nu\bar{\nu}gg$ events are generated and simulated using the Monte Carlo method. The simulation is conducted within the CEPC software framework [9], utilizing Whizard1.95 [52] and Pythia6.4 [53] for event generation, MokkaPlus [54, 55] for detector simulation, and Arbor [14, 25] for particle flow reconstruction. Using tracks and calorimeter hits as inputs, Arbor groups calorimeter hits into clusters based on the shower’s tree-like topology. With the high granularity of the AURORA calorimeter, Arbor can efficiently separate nearby showers.
For future studies, we also plan to perform comparative analysis with different generators, such as Pythia8 [56], MadGraph [57], and Herwig [58], to quantify the impact of different hadronization models on advanced reconstruction algorithms, which is crucial for the relevant uncertainty control.
A.3 Particle mapping
The visible-reconstructed particle mapping is constructed using the truth links recorded in the simulation sample. These truth links connect each tracker/calorimeter hit with the particles that excite the hit. The hits are also associated with reconstructed particles through the reconstruction procedure. The mapping is thus established via hits. By analyzing the mapping, all reconstructed particles are labeled into 15 types:
  • 1.
    10 types of 1-1 correspondence (in shades of green in Fig. 2): $e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$, $\gamma$, $K_{L}^{0}$, $n$, $\bar{n}$, and others such as $K_{S}^{0}$ and $\Lambda$.
  • 2.
    Charged particles with no track (in light green in Fig. 2): Since a track is an indispensable feature that provides more precise measurement and key information for identifying charged particles, the absence of a track will undoubtedly affect the identification of charged particles. We therefore classify this case as a separate category.
  • 3.
    Fake particles from charged shower fragments (in orange in Fig. 2).
  • 4.
    Fake particles from neutral shower fragments (in yellow in Fig. 2).
  • 5.
    Multi-track: more than one charged reconstructed particle (track) is mapped to one visible particle. Considering that the proportion of this type is less than 0.03%, we omit this case in the mapping schematic diagram for simplicity.
  • 6.
    In Arbor reconstruction, a small fraction ($<0.1\%$) of reconstructed particles cannot be mapped to any visible particle. These particles are virtually reconstructed through Arbor’s supplementary “energy flow” strategy to balance energy during track and cluster matching. Since this is an algorithm-dependent case, it is also omitted in the mapping schematic diagram.
A.4 ParT
ParT [32] is used to classify reconstructed particles through an encoder-only Transformer architecture. This model processes both individual particle features and pairwise particle interaction features. The particle features are first embedded with a 3-layer multi-layer perceptron (MLP) and then passed through 10 layers of multi-head self-attention, where the pairwise interaction features introduce a pre-softmax bias to the attention mechanism. The final particle embeddings, obtained from the last attention layer, are further processed by three fully connected layers, each with 128 neurons and a dropout rate of 0.1. A final linear layer projects the embeddings into a space with dimensions corresponding to the number of classes, generating a set of scores that represent the classification probabilities for each particle.
A.4.1 Training setup
Each reconstructed particle is characterized by 55 observable variables at the particle, cluster, and track levels, as listed in Table 2. These variables serve as input features for ParT, which outputs the likelihoods for different types of reconstructed particles. The complete sample of 1 million reconstructed $ZH\to\nu\bar{\nu}gg$ events are split into three independent sets in a ratio of 6:2:2 for training, validation, and evaluation of the ParT model. A ParT model with approximately 2.2 million parameters is constructed and trained for 30 epochs.
Table 2: Input variables of ParT.
Object level Observable variables
Reconstructed particle 4-momentum ($E,p_{x},p_{y},p_{z}$)
Direction ($\theta,\phi$)
Number of tracks and clusters
Track Number of hits
Endpoint position
3-momentum ($|\vec{p}|$, $p_{x}$, $p_{y}$, $p_{z}$, $p_{T}$)
${\rm d}E/{\rm d}x$ (mean of 5–85% truncation and quartiles)
Cluster Number of hits
Energy
Position of shower starting point
Position of center of gravity
Fractal dimension [59]
Second moment ($M_{2}$)
Distance between ECAL inner surface and shower starting point
Distance between ECAL inner surface and center of gravity
Distance between ECAL inner surface and the innermost hit
Distance between ECAL inner surface and the outermost hit
Maximum distance between cluster hits and the track helix (for charged particles)
Maximum distance between cluster hits to the axis from the innermost hit to the center of gravity
Average distance between cluster hits to the axis from the innermost hit to the center of gravity
Hit time spectrum (the fastest time and quintiles)
Closest charged cluster Minimum distance between cluster hits of each other
Number of hits
Energy
Ratio of $E_{\rm cluster}$ to $p_{\rm track}$
A.4.2 Fake particle suppression
Fake particles are those extra reconstructed neutral particles arising from shower fragments of charged particles, denoted by $f^{\pm}$. Fig. 7 compares the output $f^{\pm}$-likelihoods of different types of reconstructed neutral particles. Identified $f^{\pm}$ is defined as those reconstructed neutral ones with $f^{\pm}$-likelihood larger than some threshold. By excluding the identified $f^{\pm}$, the optimal BMR of 2.75% is derived when $f^{\pm}$-likelihood $>$ 0.85.
Figure 7: Likelihoods of fake particles from charged shower fragments ($f^{\pm}$) of different types of reconstructed neutral particles. $f^{0}$ represents fake particles from neutral shower fragments. The dashed line corresponds to the optimal working point to minimize the BMR by excluding the identified $f^{\pm}$.